OpenStudy (anonymous):

Which expression is the simplified form of √ 108x^3y^8

OpenStudy (anonymous):

\[\sqrt{108x ^{3}y ^{8}}\]

OpenStudy (anonymous):

Is it 6xy^4 √ 6x???

OpenStudy (anonymous):

\[6xy^4\sqrt{6x}\]

OpenStudy (anonymous):

\[6xy^4\sqrt{3x}\]

OpenStudy (anonymous):

Thanks:D